OpenStudy (anonymous):

solve cos^4x-1=0 for all real values of x

OpenStudy (anonymous):

first add 1 to both sides

OpenStudy (anonymous):

then you hav (cos(x))^4 = 1

OpenStudy (anonymous):

take the 4th root of each side

OpenStudy (anonymous):

don't i have to factor the cos^4x or something O_o

OpenStudy (anonymous):

cos(x) = +-1

OpenStudy (anonymous):

after that do i have to find the degrees of it? that pertains to the unit circle?

OpenStudy (anonymous):

i dont think you would need to. once you do the root of each side take the arccos(x) x=arcos(1) and arcos(-1)

OpenStudy (anonymous):

so thats the real value of x O_o\

OpenStudy (anonymous):

yeh it should be 0˚ and 180˚

OpenStudy (anonymous):

thats how i would do it

OpenStudy (anonymous):

thank you so much :D!!!!

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

so just 0 and 180 right for the real values?

OpenStudy (anonymous):

yeh because cos(x) has to equal 1 because there is nothing you can raise to the 4th power to =1 except 1 and -1. and 0 and 180 are the only things that give you 1 or -1

OpenStudy (anonymous):

I can't thank you enough WHY CANT YOU BE MORE TEACHER LOL

OpenStudy (anonymous):

remember though, thats 0 and 180 degrees. if it asks for radians make sure to convert them

OpenStudy (anonymous):

and your welcome

OpenStudy (anonymous):

i have one more

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

uhmmm...... 4cos^2x-3=0 for 0<x<2pi

OpenStudy (anonymous):

ok for this one first add 3 to both sides 4(cosx)^4=3

OpenStudy (anonymous):

then divide both sides by 4 cos(x)^4=3/4

OpenStudy (anonymous):

then take the 4th root of 3/4

OpenStudy (anonymous):

shouldnt it be a 2 instead of a 4?

OpenStudy (anonymous):

im sorry yes

OpenStudy (anonymous):

OHHHH! i gotchu, so you square root the uhmmm... 3/4 right?

OpenStudy (anonymous):

yes, and that will give you a positive and negative value

OpenStudy (anonymous):

so it would be cos^2x= 3/4 right? then square root that :O?

OpenStudy (anonymous):

yes, and that will give you a positive and negative value. then just take the arccos of those values and that should be it

OpenStudy (anonymous):

soo it would be 3/8?

OpenStudy (anonymous):

no actually what you end up with is when you take a square root its: sqrt(3/4) = (3/4)^(1/2) = (3^.5)/(4^.5) = (3^.5)/2 or sqrt(3)/2

OpenStudy (anonymous):

i have to square root 1/2 to get that O_o

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

no you should have to sqrt(1/2)

OpenStudy (anonymous):

shouldn't*

OpenStudy (anonymous):

oh okay well thank you!

OpenStudy (anonymous):

how bout uhmm.....2x^2+3x-2=0 0 degrees <x < 360degrees

OpenStudy (anonymous):

sqrt(3/4) = sqrt(3)/2 so youll have cos(x) = +-sqrt(3)/2 then just take the arccos

OpenStudy (anonymous):

ok for 2x^2+3x-2=0 0 degrees <x < 360degrees do you know how to factor

OpenStudy (anonymous):

i forgot to

OpenStudy (anonymous):

one of them has -3 and 1 in it right

OpenStudy (anonymous):

well its a bit different since you have a 2 in front of the x^2

OpenStudy (anonymous):

where do i put the 2 then

OpenStudy (anonymous):

Well normally you would have it broken into (x+somehting)(x+something) where something can be positive or negative. since you have a 2 in front itll be in the form (2x+something)(x+something)

OpenStudy (anonymous):

its a little tricky to actually figure out what those somethings are

OpenStudy (anonymous):

but from your equation 2x^2 +3x -2 you know that something*somehting has to give you -2 right

OpenStudy (anonymous):

the only things that do that are -1 and 2 or -2 and one so you can just try those in the equation and see what works it ends up being (2x-1)(x+2)

OpenStudy (anonymous):

then set 2x-1= 0 and x+2 = 0 and solve for x

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

x= 1/2 and -2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how aboutttt